# From Text to Trajectory: Exploring Complex Constraint Representation and Decomposition in Safe Reinforcement Learning

### v1

@@ 3 Preliminaries
Problem formulation. Trajectory-level constraint problem can be formed as the Constrained Non-Markov Decision Process (CNMDP) [31, 50], and it can be defined by the tuple <$S,A,T,R,\gamma,C,Y,\tau^{*}$>. Here $S$ represents the set of states, $A$ represents the set of actions, $T$ represents the state transition function, $R$ represents the reward function, and $\gamma\in(0,1)$ represents the discount factor. In addition, $Y$ represents the set of trajectory-level textual constraints (e.g., “You have 10 HP, you will lose 3 HP every time you touch the lava, don’t die.”), which describes the constraint that the agent needs to obey across the entire trajectory. $C$ represents the cost function determined by $y\in Y$. $\tau^{*}$ represents the set of historical trajectories.
@@ 4.1 Text-Trajectory Alignment Component
$\operatorname{sim}_{T}(\tau,y)=\exp(\alpha)*\frac{H_{T}\cdot L^{\top}}{||H_{T}||\text{ }||L||},\text{ }\operatorname{sim}_{T}(y,\tau)=\exp(\alpha)*\frac{L^{\top}\cdot H_{T}}{||L||\text{ }||H_{T}||},$ (4)
@@ 4.2 Cost Assignment Component
Here we regard the text representation as the $query$, each time step’s representation in the trajectory as the $key$, and compute the attention score $e_{t}$ based on the cosine similarity metric. After that, we use the sigmoid function to make sure the score falls within the range of 0 to 1. Each attention score $e_{t}$ quantifies the degree of influence of the state-action pair $(s_{t},a_{t})$ on violating the textual constraint. Then we obtain an influence-based representation $H_{t}^{*}=e_{t}H_{t}$. To predict the cost $\hat{c}(s_{t},a_{t},y,\tau_{t})$, we incorporate a feed-forward layer called the cost assignment layer $F^{c}$ and output the predicted non-violation single step cost as:
@@ 6.4 Further Results
[…]
@@ A.4 Policy Training Detail
16:     Sample batch of $N$ transitions from buffer.
@@ B.2 Case Study of Cost Assignment (CA) Component
The (a) constraint is mathematical textual constraint: “You only have 20 HP. Lava and grass are dangerous, they will make you lose 3 and 2 HP, respectively. However, water can regenerate 1 HP. Please don’t die.”. This constraint describes the two dangerous entities lava and grass, and the beneficial entity water. From the third-row heat map, we can observe that our cost assignment component assigns a high cost to the action that steps on lava or grass, with the cost increasing as the agent approaches the constraint-violating situation. Not only that, the CA component also recognizes the different levels of danger posed by lava and grass. Since stepping on lava will deduct 3 HP while stepping on grass will deduct 2 HP, the CA component assigns a larger cost value at time step $13-14$ compared to the cost value at time step $3-5$.
@@
The (c) constraint is sequential textual constraint: “After you touch lava, don’t step on grass!”. Our CA component captures two relevant entities: lava and grass, and understands the sequential relationship between entities. When the agent first steps onto the grass, the text-trajectory alignment component determines that this action does not violate the textual constraint. However, after the agent steps onto lava and then steps onto the grass for the second time, the text-trajectory alignment component detects the cosine similarity $\operatorname{sim}(\tau,y)$ greater than the threshold $\beta$, thereby violating the textual constraint. And CA component captures that the key trigger condition for violating constraint is stepping onto the lava, therefore assigning a relatively larger cost to such actions at time step $7-9$. The cost assignment component also assigns relatively small costs to some safe actions that are stepping on the floor, such as steps $13-15$. This is because it detects that the agent, although hasn’t stepped on grass yet, is trending towards approaching grass, which is a hazardous trend, thus providing a series of gradually increasing and small costs. This demonstrates that our component not only monitors key actions that lead to constraint violations but also monitors the hazardous trend and nudging the agent to choose relatively safer paths.

### v3

@@ 3 Preliminaries
Problem formulation. Trajectory-level constraint problem can be formed as the Constrained Non-Markov Decision Process (CNMDP) [31, 50], and it can be defined by the tuple $S,A,T,R,\gamma,C,Y,\tau^{*}$>. Here $S$ represents the set of states, $A$ represents the set of actions, $T$ represents the state transition function, $R$ represents the reward function, and $\gamma\in(0,1)$ represents the discount factor. In addition, $Y$ represents the set of trajectory-level textual constraints (e.g., “You have 10 HP, you will lose 3 HP every time you touch the lava, don’t die.”), which describes the constraint that the agent needs to obey across the entire trajectory. $C$ represents the cost function determined by $y\in Y$. $\tau^{*}$ represents the set of historical trajectories.
@@ 4.1 Text-Trajectory Alignment Component
[…]
@@ 4.2 Cost Assignment Component
Here we regard the text representation as the queryquery, each time step’s representation in the trajectory as the keykey, and compute the attention score $e_{t}$ based on the cosine similarity metric. After that, we use the sigmoid function to make sure the score falls within the range of 0 to 1. Each attention score $e_{t}$ quantifies the degree of influence of the state-action pair $(s_{t},a_{t})$ on violating the textual constraint. Then we obtain an influence-based representation $H_{t}^{*}=e_{t}H_{t}$. To predict the cost $\hat{c}(s_{t},a_{t},y,\tau_{t})$, we incorporate a feed-forward layer called the cost assignment layer $F^{c}$ and output the predicted non-violation single step cost as:
@@ 6.4 Further Results
Figure 6: Results of Pareto frontiers. We compare the performance of 200 policies trained using cost prediction (CP) and 200 policies trained with ground-truth cost (GC). The $\bigstar\bigstar$ symbol represents the policy on the Pareto frontier. And we connect the Pareto-optimal policies with a curve.
@@ A.4 Policy Training Detail
16:  Sample batch of NN transitions from buffer.
@@ B.2 Case Study of Cost Assignment (CA) Component
The (a) constraint is mathematical textual constraint: “You only have 20 HP. Lava and grass are dangerous, they will make you lose 3 and 2 HP, respectively. However, water can regenerate 1 HP. Please don’t die.”. This constraint describes the two dangerous entities lava and grass, and the beneficial entity water. From the third-row heat map, we can observe that our cost assignment component assigns a high cost to the action that steps on lava or grass, with the cost increasing as the agent approaches the constraint-violating situation. Not only that, the CA component also recognizes the different levels of danger posed by lava and grass. Since stepping on lava will deduct 3 HP while stepping on grass will deduct 2 HP, the CA component assigns a larger cost value at time step 131413-14 compared to the cost value at time step 353-5.
@@
The (c) constraint is sequential textual constraint: “After you touch lava, don’t step on grass!”. Our CA component captures two relevant entities: lava and grass, and understands the sequential relationship between entities. When the agent first steps onto the grass, the text-trajectory alignment component determines that this action does not violate the textual constraint. However, after the agent steps onto lava and then steps onto the grass for the second time, the text-trajectory alignment component detects the cosine similarity $\operatorname{sim}(\tau,y)$ greater than the threshold $\beta$, thereby violating the textual constraint. And CA component captures that the key trigger condition for violating constraint is stepping onto the lava, therefore assigning a relatively larger cost to such actions at time step 797-9. The cost assignment component also assigns relatively small costs to some safe actions that are stepping on the floor, such as steps 131513-15. This is because it detects that the agent, although hasn’t stepped on grass yet, is trending towards approaching grass, which is a hazardous trend, thus providing a series of gradually increasing and small costs. This demonstrates that our component not only monitors key actions that lead to constraint violations but also monitors the hazardous trend and nudging the agent to choose relatively safer paths.